Please explain and answer these two homework questions:
1) If r(x)= -2x^2-3x-3, find r(-4)
2) -3{((-4-1)-(-6^0-2)-1)-1} + or - sqroot of 9
Thanks

Answer by Boreal(15235)

r(x)= -2x^2-3x-3, find r(-4)
r(-4)=-2(-4)^2-3(-4)-3=-32+12=3=-23
-3{((-4-1)-(-6^0-2)-1)-1} + or - sqroot of 9
-3((-5)-(-1-2)-1)-1) +/-3
-3(-4)+/- 3
12+/-3
9, 15